# Machine learning a fixed point action for SU(3) gauge theory with a gauge equivariant convolutional neural network

### v1

@@ IV.3 Training
The data used to train and evaluate the network are SU(3) gauge ensembles on volumes $4^{4}$, $6^{4}$, and $8^{4}$ with the Wilson gauge action and bare gauge couplings $\beta_{\rm wil}$ ranging from 5.0 to 100.0, with more dense spacing in $\beta_{\rm wil}$ at the stronger coupling end. Each member of these ensembles represents a possible coarse configuration $V$ in Eq. (3), the minimization procedure begins with a random starting fine configuration $U$ and a parametrization of $\mathcal{A}^{\rm FP}[U]$ appropriate for smooth gauge links. Here, we use the APE444 parametrization (see Appendix C.2 for details). Minimizing $\mathcal{A}^{\rm FP}[U]+T[U,V]$ by adaptively updating of the links $U$ produces sets of fine configurations with matching volumes $8^{4},12^{4}$ and $16^{4}$. Each ensemble consists of 200 saved configurations equally spaced from Markov chains of length $10^{6}$, the ensembles are split into $80\%$ training, $10\%$ validation, and $10\%$ test data.
@@ V.1 Architecture search
The flexibility of the L-CNN architecture allows for a large variation of the network hyperparameters, namely the number of bilinear convolution layers, the number of channels, and the kernel size for convolutions. To gain some insight as to the optimal choices for these hyperparameters, we train a large set of models on the same data set, gauge ensembles with lattice volume $4^{4}$ generated with the Wilson gauge action and bare coupling $\beta_{\rm wil}$ from $5.0$ to $10.0$, for which minimization was first done to find the corresponding values of the FP action and derivatives. In the L-CNN models, we use the local Wilson action density as the prefactor $\mathcal{A}_{x}^{\mathrm{pre}}[V]$. Details about the various architectures are shown in Table 1, where we list the number of bilinear convolutional layers, and their associated kernel sizes and output channels. We also provide the number of trainable parameters. As detailed at the end of Section IV.1, the number of parameters for each bilinear convolutional grows quickly with the number of channels, the kernel size, and the number of dimensions. For each unique architecture of the thirteen listed in the table, we use both Eqs. (25) and (26) and train each architecture five times using random initial weights. This amounts to a total of 130 unique models.
@@ V.4 Finetuning for different lattice sizes
$4^{4}$ $\mathbf{7.63\times 10^{-2}}$ $8.19\times 10^{-2}$ $8.22\times 10^{-2}$
$6^{4}$ $\mathbf{7.39\times 10^{-2}}$ $7.93\times 10^{-2}$ $7.96\times 10^{-2}$
$8^{4}$ $\mathbf{7.36\times 10^{-2}}$ $7.91\times 10^{-2}$ $7.93\times 10^{-2}$
@@ VI Conclusions and outlook
In the challenge of pushing lattice QCD calculations to ever higher accuracy one faces the imminent problems of critical slowing down and topological freezing. In this paper, we propose to overcome these obstacles by using highly improved gauge actions, such that simulations at coarse lattice spacing are possible, where critical slowing down and topological freezing are avoided, while keeping lattice artifacts under sufficient control to take a reliable continuum limit. In order to do so, we follow the FP action approach [18] based on the properties of RG transformations. FP actions are lattice discretizations which have no lattice artifacts at the classical level. They are also expected to have suppressed lattice artifacts at the quantum level. Parametrizing the complicated FP actions has been a major challenge in the past [25] and in this paper we address it by employing a gauge equivariant convolutional neural network (L-CNN) [30] and ML techniques.
@@
The main observation in this paper is that trained L-CNN models are able to achieve much higher accuracy than previous parametrizations of the FP action over a larger range of lattice spacings and corresponding gauge field fluctuations. This is not surprising, given the flexibility and large number of parameters of the neural network models. It is particularly encouraging that the L-CNN accuracy varies little in the range of coarsest lattice spacings as shown in Fig. 9. The baseline FP parametrization APE431 was previously used in Monte Carlo studies of the deconfinement phase transition, the static quark-antiquark potential, and the glueball mass spectrum [25], with the promising result that the parametrized FP action had very small lattice artifacts, if any, in these physical observables up to lattice spacings as coarse as $a\sim 0.33$ fm, at the level of accuracy feasible at that time.
@@
The ultimate test of the L-CNN parametrization of the FP action will be its performance in actual Monte Carlo simulations. In the course of this project, we gained valuable experience in determining derivatives of the FP action with respect to gauge links through backpropagation. This method opens up interesting possibilities for simulation methods based on derivatives such as HMC algorithms [49] or Langevin dynamics [50, 51, 52]. Both simulation strategies use the variation of the action with respect to the gauge fields. Since the derivatives can be calculated efficiently within the L-CNN architecture and constitute the major part of the data used for the learning of the L-CNN, this aids the feasibility of large-scale simulations.
@@
A more ambitious and difficult goal — the holy grail of RG improvement — is to find the exact RG trajectory as in Fig. 1, for which cutoff artifacts are eliminated completely. Such a quantum perfect action (or RG action) would in principle allow to extract continuum physics from simulations at one (coarse) lattice spacing. A procedure to determine the RG action using gauge-field derivatives is outlined in [18]. Constructing and parametrizing the RG action is another challenge which can be embraced using the L-CNN and ML techniques. The FP action constructed in this paper is a crucial and necessary first step in that direction.
@@ Acknowledgments
The authors wish to thank Anna Hasenfratz, Jim Hetrick, Gurtej Kanwar and Uwe-Jens Wiese for valuable discussions. This material is based upon work supported by the US National Science Foundation under Grant No. 2014150. KH and DM wish to thank the AEC and ITP at the University of Bern for their support. KH, AI and UW thank ECT* and the ExtreMe Matter Institute EMMI at GSI, Darmstadt, for support in the framework of the ECT*/EMMI Workshop “Machine learning in lattice field theory and beyond” during which this work has been developed. AI and DM acknowledge funding from the Austrian Science Fund (FWF) projects P 32446, P 34455 and P 34764. The computational results presented have been achieved in part using the Vienna Scientific Cluster (VSC) and computing resources at the University of Bern.
@@ Appendix D Renormalization group transformation
$\mathsf{W}_{n,\mu}=\sum_{m=0}^{3}\sum_{n^{\prime}}s_{m}W^{(m)}_{n,n^{\prime}}U_{n^{\prime},\mu}W^{(m)}_{n^{\prime}+\hat{\mu},n+\hat{\mu}}.$ (67)

### v2

@@ IV.3 Training
The data used to train and evaluate the network are SU(3) gauge ensembles on volumes $4^{4}$, $6^{4}$, and $8^{4}$ with the Wilson gauge action and bare gauge couplings $\beta_{\rm wil}$ ranging from 5.0 to 100.0, with more dense spacing in $\beta_{\rm wil}$ at the stronger coupling end. Each member of these ensembles represents a possible coarse configuration $V$ in Eq. (3), the minimization procedure begins with a random starting fine configuration $U$ and a parametrization of $\mathcal{A}^{\rm FP}[U]$ appropriate for smooth gauge links. Here, we use the APE444 parametrization (see Appendix C.2 for details). Minimizing $\mathcal{A}^{\rm FP}[U]+T[U,V]$ by adaptively updating of the links $U$ produces sets of fine configurations with matching volumes $8^{4},12^{4}$ and $16^{4}$. Each ensemble consists of 200 saved configurations equally spaced from Markov chains of length $10^{6}$, and the ensembles are split into $80\%$ training, $10\%$ validation, and $10\%$ test data.
@@ V.1 Architecture search
The flexibility of the L-CNN architecture allows for a large variation of the network hyperparameters, namely the number of bilinear convolution layers, the number of channels, and the kernel size for convolutions. To gain some insight as to the optimal choices for these hyperparameters, we train a large set of models on the same data set, gauge ensembles with lattice volume $4^{4}$ generated with the Wilson gauge action and bare coupling $\beta_{\rm wil}$ from $5.0$ to $10.0$, for which minimization was first done to find the corresponding values of the FP action and derivatives. In the L-CNN models, we use the local Wilson action density as the prefactor $\mathcal{A}_{x}^{\mathrm{pre}}[V]$. Details about the various architectures are shown in Table 1, where we list the number of bilinear convolutional layers, and their associated kernel sizes and output channels. We also provide the number of trainable parameters. As detailed at the end of Section IV.1, the number of parameters for each bilinear convolution grows quickly with the number of channels, the kernel size, and the number of dimensions. For each unique architecture of the thirteen listed in the table, we use both Eqs. (25) and (26) and train each architecture five times using random initial weights. This amounts to a total of 130 unique models.
@@ V.4 Finetuning for different lattice sizes
$4^{4}$ $7.63\times 10^{-2}$ $8.19\times 10^{-2}$ $8.22\times 10^{-2}$
$6^{4}$ $7.39\times 10^{-2}$ $7.93\times 10^{-2}$ $7.96\times 10^{-2}$
$8^{4}$ $\mathbf{7.36\times 10^{-2}}$ $\mathbf{7.91\times 10^{-2}}$ $\mathbf{7.93\times 10^{-2}}$
@@ VI Conclusions and outlook
In the challenge of pushing lattice QCD calculations to ever higher accuracy one faces the imminent problems of critical slowing down and topological freezing. In this paper, we propose to overcome these obstacles by using highly improved gauge actions, such that simulations at coarse lattice spacing are possible, where critical slowing down and topological freezing are avoided, while keeping lattice artifacts under sufficient control to take a reliable continuum limit. In order to do so, we follow the FP action approach [18] based on the properties of RG transformations. FP actions are lattice discretizations which have no lattice artifacts at the classical level. They are also expected to have suppressed lattice artifacts at the quantum level. Parametrizing the complicated FP actions has been a major challenge in the past [25] and in this paper we address it by employing a gauge equivariant convolutional neural network (L-CNN) [30] and ML techniques. Studying the quality and improvement achieved with the new parametrization is the first conceptual step in the construction of highly improved RG actions.
@@
The main observation in this paper is that trained L-CNN models are able to achieve much higher accuracy than previous parametrizations of the FP action over a larger range of lattice spacings and corresponding gauge field fluctuations. This is not surprising, given the flexibility and large number of parameters of the neural network models. It is particularly encouraging that the L-CNN accuracy varies little in the range of coarsest lattice spacings as shown in Fig. 9. The baseline FP parametrization APE431 was previously used in Monte Carlo studies of the deconfinement phase transition, the static quark-antiquark potential, and the glueball mass spectrum [25], with the promising result that the parametrized FP action had very small lattice artifacts in these physical observables up to lattice spacings as coarse as $a\sim 0.33$ fm, at the level of accuracy feasible at that time. In the same spirit, the ultimate test of the L-CNN parametrization of the FP action will of course be its performance in actual Monte Carlo simulations and conducting state-of-the-art scaling tests on coarse lattices is therefore the crucial next step in our attempt to construct highly improved RG actions.
@@
In the course of this project, we gained valuable experience in determining derivatives of the FP action with respect to gauge links through backpropagation. This method opens up interesting possibilities for simulation methods based on derivatives such as HMC algorithms [49] or Langevin dynamics [50, 51, 52]. Both simulation strategies use the variation of the action with respect to the gauge fields. Since the derivatives can be calculated efficiently within the L-CNN architecture and constitute the major part of the data used for the learning of the L-CNN, this aids the feasibility of large-scale simulations.
@@
A more ambitious and difficult goal — the holy grail of RG improvement — is to find the exact RG trajectory as in Fig. 1, for which cutoff artifacts are eliminated completely. Such a quantum perfect action (or RG action) would in principle allow one to extract continuum physics from simulations at one (coarse) lattice spacing. A procedure to determine the RG action using field derivatives is outlined in [18]. Constructing and parametrizing the RG action is another challenge which can be embraced using the L-CNN and ML techniques. The FP action constructed in this paper is a crucial and necessary first step in that direction.
@@ Acknowledgments
The authors wish to thank Anna Hasenfratz, Jim Hetrick, Gurtej Kanwar and Uwe-Jens Wiese for valuable discussions. This material is based upon work supported by the US National Science Foundation under Grant No. 2014150. KH and DM wish to thank the AEC and ITP at the University of Bern for their support. KH, AI and UW thank ECT* and the ExtreMe Matter Institute EMMI at GSI, Darmstadt, for support in the framework of the ECT*/EMMI Workshop “Machine learning in lattice field theory and beyond” during which this work has been developed. AI and DM acknowledge funding from the Austrian Science Fund (FWF) projects P 32446, P 34455 and P 34764. UW acknowledges funding from the Swiss National Science Foundation (SNSF) project No. 200020_208222. The computational results presented have been achieved in part using the Vienna Scientific Cluster (VSC) and computing resources at the University of Bern.
@@ Appendix D Renormalization group transformation
$\mathsf{W}_{n,\mu}=\sum_{m=0}^{3}\sum_{n^{\prime}}s_{m}W^{(m)}_{n,n^{\prime}}U_{n^{\prime},\mu}W^{(m)}_{n^{\prime}+\hat{\mu},n+\hat{\mu}},$ (67)